OpenStudy (butterflydreamer):

help please :) - BINOMIAL THEOREM CC's question below:

OpenStudy (butterflydreamer):

Prove: \[1 + \left(\begin{matrix}10 \\ 2\end{matrix}\right)3^{2} + \left(\begin{matrix}10 \\ 4\end{matrix}\right)3^{4} + \left(\begin{matrix}10 \\ 6\end{matrix}\right)3^{6} + \left(\begin{matrix}10 \\ 8\end{matrix}\right)3^{8} + 3^{10} = 2^9 (2^{10} +1)\]

OpenStudy (butterflydreamer):

I'm considering: \[(1+x)^{10} = \left(\begin{matrix}10 \\ 0\end{matrix}\right) +\left(\begin{matrix}10 \\ 1\end{matrix}\right)x + \left(\begin{matrix}10 \\ 2\end{matrix}\right)x^2 + \left(\begin{matrix}10 \\ 3\end{matrix}\right)x^3 + ... + x^{10} \] But i have no idea where to go from here. If i look at the RHS, then i came up with something....?: \[2^9 ( 2^{10} + 1) \rightarrow 2^{n-1} (2^{n} +1) \] But i'm not how to get there :/ I also noticed on the LHS, the things are all even. Idk -.-

OpenStudy (amistre64):

``` \left(\begin{matrix}10 \\ 2\end{matrix}\right) ``` is better writen as ``` \binom{10}{2} ``` \[\binom{10}{2}\]

OpenStudy (butterflydreamer):

oh, thanks for the tip LOL

OpenStudy (amistre64):

(x-1)^(10) = cx^10 -cx^9 +cx^8 -cx^7.. right?

OpenStudy (amistre64):

what is: (x+1)^(10) + (x-1)^(10)

OpenStudy (amistre64):

\[2 (x^{10}+45 x^8+210 x^6+210 x^4+45 x^2+1)\]

OpenStudy (butterflydreamer):

ohhh true true . That makes sense! So that's how they got the even parts only.

OpenStudy (amistre64):

im an idiot, so i just use whatever comes to mind :)

OpenStudy (amistre64):

oh im not sure how THEY did it .. but this might be a start

OpenStudy (butterflydreamer):

yeess. LOL i can't believe i expanded wrong xD oh gosh. Okay sooo, what should we do next?

OpenStudy (amistre64):

dunno offhand, i was just trying to get your thought to get rid of the odd parts

OpenStudy (amistre64):

since LHS and RHS are finite numbers, cant we just let x=3 and work the math to prove?

OpenStudy (amistre64):

(x+1)^(10) + (x-1)^(10) 4^(10) + 2^(10) 2^(10) 2^10 + 2^(10) 2^(10) [2^(10)+1]

OpenStudy (amistre64):

and divide by 2, since we are twice as heavy

OpenStudy (amistre64):

or something like that

OpenStudy (butterflydreamer):

.... oh... o.o

OpenStudy (butterflydreamer):

can we solve the question like that though? LOL. I've been trying to sub all kinds of weird numbers everywhere and trying to integrate things -.- I've been getting nowhere tbh LOl

OpenStudy (amistre64):

direct solutions tend to work out great to me ... \[1 + \left(\begin{matrix}10 \\ 2\end{matrix}\right)x^{2} + \left(\begin{matrix}10 \\ 4\end{matrix}\right)x^{4} + \left(\begin{matrix}10 \\ 6\end{matrix}\right)x^{6} + \left(\begin{matrix}10 \\ 8\end{matrix}\right)x^{8} + x^{10} =\frac{(x+1)^{10}+(x-1)^{10}}{2}\]

OpenStudy (amistre64):

when x=3, ive already worked the numerator to 2^(10) [2^(10)+1] right?

OpenStudy (butterflydreamer):

yes

OpenStudy (amistre64):

then it is proofed

OpenStudy (butterflydreamer):

hmm okay then. Thank you! It's just that for the prev. questions i've been working on, i've had to use a range of differentiation/integration and substitution methods so i thought this question would've involved similar steps but thanks again! Your method was way simpler.

OpenStudy (amistre64):

\[1 + \left(\begin{matrix}10 \\ 2\end{matrix}\right)3^{2} + \left(\begin{matrix}10 \\ 4\end{matrix}\right)3^{4} + \left(\begin{matrix}10 \\ 6\end{matrix}\right)3^{6} + \left(\begin{matrix}10 \\ 8\end{matrix}\right)3^{8} + 3^{10}\] \[\frac{(3+1)^{10}+(3-1)^{10}}{2}\] \[\frac{4^{10}+2^{10}}{2}\] \[\frac{2^{10}(2^{10}+1)}{2}\] \[{2^{9}(2^{10}+1)}\]

OpenStudy (amistre64):

well, it was your way, i just built on it :)

OpenStudy (butterflydreamer):

aahh okay! Thanks again :D!